If cot x = tan nx , find the general solution for x.

Grade:11

1 Answers

arpit singh
6 years ago
If cotx=tannxtan(π/2-x)=tannxSo tan nx=tan(π/2-x) nx=nπ+π/2-x nx+x=nπ+π/2 x(n+1)=π(n+1/2) x=π/n+1(n+1/2)